# JUREX-4E: Juridical Expert-Annotated Four-Element Knowledge Base for Legal Reasoning

### v1

@@ Abstract
The Four-Element Theory is a fundamental framework in criminal law, defining the constitution of crime through four dimensions: Subject, Object, Subjective aspect, and Objective aspect. This theory is widely referenced in legal reasoning, and many Large Language Models (LLMs) attempt to incorporate it when handling legal tasks. However, current approaches rely on LLMs’ internal knowledge to incorporate this theory, often lacking completeness and representativeness. To address this limitation, we introduce JUREX-4E, an expert-annotated knowledge base covering 155 criminal charges. It is structured through a progressive hierarchical annotation framework that prioritizes legal source validity and employs diverse legal interpretation methods to ensure comprehensiveness and authority. We evaluate JUREX-4E on the Similar Charge Distinction task and apply it to Legal Case Retrieval, demonstrating its effectiveness in improving LLM performance. Experimental results validate the high quality of JUREX-4E and its substantial impact on downstream legal tasks, underscoring its potential for advancing legal AI applications. Code: https://github.com/THUlawtech/JUREX
@@ 1 Introduction
In legal AI tasks, enhancing the accuracy and interpretability of Large Language Models (LLMs) in the legal domain often requires the incorporation of legal theories as a support Jiang and Yang (2023); Servantez et al. (2024); Yuan et al. (2024); Deng et al. (2023). One important theory is the Four-Element Theory of Crime Constitution in Chinese criminal law Liang (2017). This theory deconstructs criminal conduct into four elements: Subject, Object, Subjective aspect, and Objective aspect, providing clear standards for judicial authorities to determine criminal behavior and helping to prevent the abuse of penal power.
@@
However, most current approaches do not provide additional knowledge but rather rely on the LLM’s internal knowledge to incorporate the Four-Element Theory. A common method is to guide LLMs in mimicking expert reasoning processes. For example, designing four separate prompts to guide the LLM outputs in the form of four elementsDeng et al. (2023).
These methods assume that the model has a solid grasp of the Four-Element Theory, which has not yet been verified. We had LLMs generate the four elements of several complicated crimes in Chinese judicial practiceOuyang et al. (1999), and then asked legal experts to score them. We found that, although LLMs can generate formally standardized and relatively accurate legal descriptions when provided with legal theoretical frameworks and references, the model still underperformed in terms of completeness and representativeness. This shortcoming could affect the accuracy and soundness of subsequent reasoning.
@@
To help LLMs better utilize the Four-Element Theory in legal tasks, we propose JUREX-4E: JURidical EXpert-annotated 4-Element knowledge base for legal reasoning. This knowledge base is annotated using a progressive hierarchy: Article → Judicial Interpretations → Guiding Cases → Academic Discourses, which is built upon the pyramid structure of legal source validity. It incorporates various legal interpretation methods, including textual, systematic, sociological, and purposive interpretations. The knowledge base covers the four elements of 155 high frequency charges, annotated by legal experts over a period of seven months. Each crime’s four elements are described in an average of 472.5 words.
To assess the quality of the annotations, we sampled several crimes for human evaluation. The expert annotations achieved an average score of 4.60 on a 5-point scale, while the LLM-generated four elements scored only 3.96, indicating that the expert annotations were of higher quality. To further evaluate the annotations objectively and comprehensively, a direct way is to judge whether different charges can be distinguished according to the four-element definition of crime constitution. Therefore, we introduced the Similar Charge Distinction task Liu et al. (2021). For each case, we provided the four elements of the candidate confused charges and combined them with the case facts as model input. The experimental results showed that injecting expert annotations helped the model better differentiate between similar charges, improving performance with a 0.65 increase in average accuracy and a 0.70 increase in average F1-score, underscoring the superior quality and reliability of expert annotations compared to those generated by the LLM.
We also applied the expert annotations in a specific legal task: Legal Case Retrieval. It is an important step in the practice of analyzing cases and making judgments, requiring the precise application of the four-element theory to compare the criminal composition of cases. We designed a simple retrieval framework guided by expert knowledge, in which the charge’s four elements was used to generate four-element descriptions for both the query and candidate cases, and then match similar cases based on their vector similarity. Experiments demonstrated that incorporating expert-annotated four elements improved retrieval performance, as the model became better at focusing on the legal features and key details.
@@
    We verify that LLMs have gaps in understanding the legal theory, highlighting the inadequacy of relying solely on LLM-driven reasoning for legal AI tasks.
@@
    We built the JUREX-4E knowledge base, which is the first to incorporate the pyramid structure of legal source validity and covers the four elements of 155 criminal charges under Chinese Criminal Law.
@@
    We demonstrated the significance of incorporating criminal composition elements in the Similar Charge Distinction task and proved the superior quality of the expert-annotated four-element knowledge base.
  4. (4)
    We applied JUREX-4E to the Legal Case Retrieval task, found that they do indeed contribute to downstream tasks.
@@ 2 Related Work
In legal AI, much work has introduced legal theories to enhance reasoning and improve model accuracy and interpretability. For example, legal syllogism prompting (LoT)Jiang and Yang (2023) teaches LLMs for legal judgment prediction by instructing legal syllogism, Chain of LogicServantez et al. (2024) guides models in reasoning about compositional rules by decomposing logical statements based on the IRAC (Issue, Rule, Application, Conclusion) paradigm. Among these, the Four-Elements Theory (FET) of Crime Constitution is a widely adopted frameworkYuan et al. (2024); Deng et al. (2023).
The Four-Element Theory is one of the most widely recognized criminal theories in Chinese judicial practice Liang (2017). It specifies four essential elements that must be satisfied to establish criminal liability: Subject, Object, Subjective aspect, and Objective aspect. For example, the four elements of the Crime of Affray can be briefly summarized as follows:
(1) Subject: Principal organizers and other active participants who have reached the age of criminal responsibility. (2) Object: Public order. (3) Objective Aspect: The act of assembling brawl, engaging in a brawl, resulting in the following consequences of serious injury. (4) Subjective Aspect: Direct intent, where the person knowingly and willfully engages in organizing or participating in the act of assembling brawl.
Before discussing the Four-Element Theory (FET), it is necessary to briefly compare it with another key theory in Chinese criminal law, the Hierarchical Theory of Crime ConstitutionZhou (2017b); Zhang (2010), and the main distinction between these theories lies in whether a hierarchical structure is considered, with ongoing debates in practiceGao (2009); Chen (2010, 2017); Zhou (2017a). We chose FET as our foundational template for following reasons: 1) its dominance in Chinese judicial practice aligns with real-world criminal judgments; (2) its clear distinction between objective aspects and subjective intent offers direct reasoning checkpoints compared to the Three-Tier Theory; (3) its four-element annotation is flexible and can be adapted to the Three-Tier Theory by prioritizing objective analysis before subjective evaluationLi (2006); Zhang (2017).
Recent approaches have leveraged the FET framework to model expert reasoning. For example, breaking down legal rules into FET-aligned components using automated planning techniques Yuan et al. (2024). Employing model-generated four-element structures as minor premises in legal judgment analysis Deng et al. (2023). While these methods have demonstrated improved performance on downstream tasks, they generally assume that the LLMs inherently understand the FET, without systematically validating this assumption. Notably, prior research on criminal charge prediction An et al. (2022) suggests that the models may misinterpret key legal concepts and may not be sensitive enough to the subtle differences in fact descriptions of confusing charges, highlighting the need to incorporate expert annotations to support LLM reasoning.
@@ 3.1 Hierarchical Legal Interpretation System
Figure 1: Hierarchical Legal Interpretation System base on legal source validity. The legal sources follow a hierarchical order of validity. Thick arrows indicate the primary level where a particular interpretive method is applied, while dashed arrows represent its supplementary use at that level.
Annotating the four elements of crime constitution is essentially a process of legal interpretation, which can be analyzed in two aspects:
@@
(1) What law is being interpreted. This involves identifying the sources of law, including statutory provisions corresponding to a specific charge, their associated judicial interpretations, case precedents, and academic discourses. In legal studies, these sources are categorized based on their legal validity into formal sources (which carry legal forces in judgments) and informal sources (which serve as references without legal forces)Pound (1925); Watson (1982); Pound (1932). Articles and judicial interpretations are considered formal sources, whereas case precedents and academic discourses are regarded as informal sources under the Chinese legal systemZhang and Zhou (2007).
@@
(2) How the law is interpreted. This pertains to legal interpretation methods, including literal interpretation, systematic interpretation, purposive interpretation, etc. These methods follow a hierarchical order in legal reasoningSutherland (1891); Kim and Division (2008); Eig Larry (2014). Legal interpretation should begin with literal interpretation (textual analysis). If the intended meaning cannot be clearly derived from the article alone, systematic interpretation and purpose interpretation should be applied first. If ambiguity remains, historical interpretation and comparative law interpretation may be used to further clarify the legal meaning. The specific definition is in AppendixB.
Based on these principles, our annotation follows a pyramid structure of Hierarchical Legal Interpretation System base on legal source validity. As shown in Figure 1, the system is divided into two parts: Legal Source and Legal Interpretation Methods. The main structure of legal source follows a hierarchical order of validity: Article → Judicial Interpretations → Guiding Cases → Academic Discourses, where various legal interpretation methods are applied across different levels. Thick arrows indicate the primary level at which a particular method is used, while thin arrows denote the cross applications.
3.2 Hierarchical Annotation Path of Legal Sources
Our Annotators are experts who have all passed the National Uniform Legal Profession Qualification Examination and are familiar with the Four-Element Theory. The entire annotation process took a total of 7 months and involved 4 rounds of annotation according to the validity of the legal source from high to low level.
The First Level: Article.
Legal elements can be seen as an interpretation and refinement of the statutory provisions corresponding to a particular crime. Using literal interpretation as the primary method, the statute is broken down based on its semantic meaning and common usage, ensuring that the interpretation does not extend beyond the possible meaning of the text: (1) linguistic analysis follows the subject-predicate-object structure of the provision. (2) To maintain consistency, terms are systematically classified and mapped(e.g: subjective aspect is classified as either intentional or negligent. )(3) Only when it is impossible to make an explicit inclusion or exclusion judgment for an element based on the rules of language use (neutral option field), other interpretation methods should be used. This initial phase takes almost 2 months.
@@ 3.2 Hierarchical Annotation Path of Legal Sources
For example, in the crime of robbery, the object "public or private property" represents the protected legal interest. The phrase "forcibly seizing public or private property through violence, coercion, or other means" describes the objective aspect. Since no subject is specified, it is assumed to involve a general subject, and the adverbs "violence" and "coercion" indicate an intentional act. Preliminarily interpret ‘violence’ in the objective aspect as ‘Use of physical force or power’, but the specific forms and subjects of violence need further clarification.
The Second Level: Judicial Interpretation.
In the 3rd and 4th months, the second level of the hierarchical annotation path focuses on refining legal elements through judicial interpretation. The primary method used for interpreting these materials is systematic interpretation. This approach examines the position of the corresponding articles within the legal system by analyzing their placement within the structure of laws, including parts, chapters, sections, articles, clauses, and subclauses, as well as their relationship to other statutes and judicial interpretations. Additionally, other interpretative methods, such as sociological interpretation and teleological interpretation, are referenced based on judicial interpretations, related statutory provisions, or bar exam questions.The goal of this level is to clarify the legislative intent by considering the contextual relevance of each provision within the broader legal framework.
For example, in the first level, the objective aspect of "violence" in the crime of robbery requires further clarification, specifically regarding whether violence must be directed exclusively at persons or could also apply to property. Article 289 of Chinese Criminal LawCongress (2017) stipulates that in cases of "smashing, looting, and robbing" committed by a group, the ringleaders shall be convicted of robbery if they destroy or seize public or private property. This provision demonstrates that violence against property can also constitute robbery under Chinese law.
The Third Level: Guiding Cases.
In the 5th to 6th month, purposive interpretation and sociological interpretation are applied to the guiding cases and landmark judgments from the Supreme Court. By examining the social significance of real-world cases, these methods bridge the subtle gap between abstract legal theory and practical cases. This approach enables dynamic adaptation and integration of empirical insights and emerging controversies within the dataset.
For example, in Criminal Trial Reference Case No.159Zou (2002), the perpetrator lured the victim into a room, locked the door, and seized 170,000 RMB intended for a transaction. The court determined that although the detention did not endanger personal safety, it was sufficient to suppress the victim’s resistance, thus constituting "violence" in in the objective aspects of robbery. Another example is the "Molestation and Theft Case" Ma (2021), where the perpetrator bound the victim, committed molestation, and stole the victim’s phone. Since the ongoing molestation reinforced coercion, it constitutes a new act of violence. Thus, the annotation includes "molestation" as an additional method.
The Fourth Level: Academic Discourses.
In the 7th month, the final stage involves academic expansion. Academic controversies are introduced by employing multiple interpretive methods such as comparative law interpretation, purposive interpretation, and sociological interpretation. These methods include inserting conflict markers at key points of controversy, highlighting the distinctions between mainstream consensus and minority theories, while providing brief annotations of their legal reasoning. This approach ensures the extensibility and academic depth of the dataset.
For example, regarding the crime of robbery, for the main view in China, Soviet Union, North Korea, and Japan explicitly holds that the violence must be severe enough to endanger the victim’s life or healthZhang (2007). But some scholars argue that any violence that can forcibly impact the victim’s body is sufficient to constitute violence in robbery, no need to endanger the victim’s life or healthYang (2010).
4 Data Distribution
Metric LLM Expert
Mean Median Mean Median

### v2

@@ Abstract
In recent years, Large Language Models (LLMs) have been widely applied to legal tasks. To enhance their understanding of legal texts and improve reasoning accuracy, a promising approach is to incorporate legal theories. One of the most widely adopted theories is the Four-Element Theory (FET), which defines the crime constitution through four elements: Subject, Object, Subjective Aspect, and Objective Aspect. While recent work has explored prompting LLMs to follow FET, our evaluation demonstrates that LLM-generated four-elements are often incomplete and less representative, limiting their effectiveness in legal reasoning. To address these issues, we present JUREX-4E, an expert-annotated four-element knowledge base covering 155 criminal charges. The annotations follow a progressive hierarchical framework grounded in legal source validity and incorporate diverse interpretive methods to ensure precision and authority. We evaluate JUREX-4E on the Similar Charge Disambiguation task and apply it to Legal Case Retrieval. Experimental results validate the high quality of JUREX-4E and its substantial impact on downstream legal tasks, underscoring its potential for advancing legal AI applications. The dataset and code are available at: https://github.com/THUlawtech/JUREX
@@ 1 Introduction
Large Language Models (LLMs) have recently demonstrated impressive performance in legal tasks such as charge prediction Yuan et al. (2024) and legal case retrieval Feng et al. (2024). In these applications, a key challenge is accurately understanding complex legal language. To address this, recent studies have introduced legal theories into LLM workflows Jiang and Yang (2023); Servantez et al. (2024); Yuan et al. (2024); Deng et al. (2023), as these theories provide structured reasoning frameworks and domain knowledge. Among these theories, the Four-Element Theory (FET) in Chinese criminal law Liang (2017) is particularly important, as it defines the legal criteria for establishing criminal liability. FET breaks down a criminal charge into four elements: Subject, Object, Subjective Aspect, and Objective Aspect, which serve as the essential criteria for determining whether a defendant’s behavior constitutes a specific crime.
@@
Most current approaches rely on the LLM’s internal knowledge to incorporate the FET. A common method is to ask LLMs to emulate expert reasoning processes. For example, designing four separate prompts to guide the LLM outputs in the form of four-elements Deng et al. (2023). This raises a critical question: Can LLMs reliably understand and apply the FET?
Figure 1: An example of LLM-generated four-elements.
To investigate this, we conducted a pilot study where we provide LLMs with legal articles and asked them to generate the four-elements for several representative charges Ouyang et al. (1999). Results show that the LLM-generated four-elements are often not accurate enough. As shown in Figure 1, in the charge of misappropriation of public funds, the LLM failed to identify the right to benefit from the use of public funds, a core part of the Object. These results suggest that LLMs lack the domain knowledge and legal reasoning precision required for reliable FET application.
@@
To help LLMs better utilize the FET in legal tasks, we construct JUREX-4E: JURidical EXpert-annotated 4-Element knowledge base for legal reasoning. The knowledge base covers 155 high-frequency criminal charges, each decomposed into Subject, Object, Subjective Aspect, and Objective Aspect. JUREX-4E is built through a four-stage Hierarchical Legal Interpretation System. In this process, legal experts refine each element by referencing sources in descending order of legal validity—Criminal Articles, Judicial Interpretations, Guiding Cases, and Academic Discourses—while applying appropriate interpretive methods at each stage. Each charge was annotated over a seven-month period, yielding knowledge-rich representations with an average annotation length of 472.5 words.
To assess the quality of JUREX-4E, we conduct a human evaluation on four independent dimensions, Precision, Completeness, Representativeness, and Standardization, grounded in legal scholarship on how criminal elements should be normatively defined and expressed in judicial contexts  Zhang (2007a). The expert-annotated four-elements achieved an average score of 4.60 on a 5-point scale, significantly outperforming the LLM-generated ones, which scored 3.96. Among the four dimensions, the largest performance gaps appeared in Completeness and Representativeness, as expert annotations provided more comprehensive legal interpretations and summarized typical application scenarios, which are often overlooked by LLMs.
To further evaluate the quality and utility of JUREX-4E, we conducted two downstream tasks: Similar Charge Disambiguation (SCD) and Legal Case Retrieval (LCR). In the SCD task Liu et al. (2021), we tested whether different charges could be more effectively distinguished by incorporating four-element knowledge. Results show that expert-annotated four-elements from JUREX-4E consistently outperformed LLM-generated counterparts across various prompting strategies and model types, improving average accuracy by 0.70% and F1-score by 0.75%. In the LCR task Li et al. (2024d), we incorporated JUREX-4E into the retrieval pipeline to guide case-level four-element generation and similarity matching, achieving better retrieval accuracy. Together, these findings validate the high quality and practical value of JUREX-4E in enhancing legal understanding and decision-making.
@@
    We demonstrate that while LLMs can assist legal reasoning to some extent, they still fall short in accurately understanding and applying the Four-Element Theory.
@@
    We construct the JUREX-4E, the first expert-annotated legal knowledge base grounded in a hierarchical legal interpretation framework based on legal source validity.
@@
    We validate the quality and effectiveness of JUREX-4E on two representative legal tasks, Similar Charge Disambiguation (SCD) and Legal Case Retrieval (LCR), where it consistently outperforms LLM-generated representations across various prompting strategies.
2 Background
The Four-Element Theory (FET) of crime constitution is a fundamental framework in Chinese criminal law Liang (2017). It provides a standardized structure to determine criminal liability through four elements: Subject, Object, Subjective Aspect, and Objective Aspect.
For example, the four-elements of Robbery can be briefly summarized as follows:
(1) Subject (the person who commits a criminal act and should bear criminal responsibility): General subjects above the age of criminal responsibility.
(2) Object (the legal interest harmed by the act): A compound object, combining both property ownership and personal rights of the victim.
(3) Subjective Aspect (the offender’s mental state regarding the harmful act): Direct intent with the purpose of unlawfully appropriating another’s property.
(4) Objective Aspect (the external facts of the criminal activity, including key actions and their outcomes): On-the-spot taking of property from an owner, custodian, or possessor through violence, coercion, or other methods.
For the legal community, FET plays a central role in doctrinal analysis and judicial reasoning. It serves as the legal basis for both legislation and adjudication, ensuring internal consistency and normative rigor in criminal law application Li (2006); Zhang (2007a). For the legal AI community, FET offers a task-agnostic and interpretable framework for modeling legal reasoning Deng et al. (2023); Yuan et al. (2024).
Compared to general reasoning templates (e.g., legal syllogism Gold (2018)) or alternative theories such as the Three-Tier System Zhou (2017); Zhang (2010), FET has become the dominant approach in China for assessing criminal liability Wang (2017). Its clearer and more interpretable decomposition of crimes into objective and subjective elements makes it particularly suitable for structured legal reasoning tasks.
@@ 3 Related Work
With the rise of open-source base LLMs, lots of legal LLMs have emerged, such as Lawyer LLaMA Huang et al. (2023), DiSC-LawLLM Yue et al. (2023), ChatLaw Cui et al. (2024), and TongyiFarui [^1] https://tongyi.aliyun.com/farui. These models are typically adapted from general-purpose LLMs via domain-specific post-training or Retrieval-Augmented Generation (RAG), incorporating legal texts like cases and laws.
Although these models achieve notable improvements on legal tasks, they still struggle with complex legal reasoning, such as charge disambiguation, legal question answering, statutory interpretation, and structured explanation generation Hu et al. (2025). LegalDiscourse shows that LLMs often fail to capture when laws apply and to whom Spangher et al. (2024), while LegalBench demonstrates that even state-of-the-art models underperform on diverse reasoning-intensive legal tasks Guha et al. (2023).
To further enhance model performance, particularly in tasks requiring complex legal reasoning, some studies draw inspiration from established legal reasoning paradigms. For example, introducing the legal syllogism for legal judgment prediction Jiang and Yang (2023); using the IRAC paradigm to guide LLMs in reasoning about compositional rules Servantez et al. (2024). Several works have drawn on the FET in the context of Chinese criminal law. For example, breaking down legal rules into FET-aligned components using automated planning techniques Yuan et al. (2024); employing model-generated FETs as minor premises in legal judgment analysis Deng et al. (2023).
While these methods have demonstrated improved performance on downstream tasks, they generally assume that the LLMs inherently understand the Four-Element Theory, without systematically validating this assumption.
4 Can LLMs Grasp Legal Theory?
To examine whether LLMs can understand and apply the Four-Element Theory (FET), we ask them to generate the four elements (FETs) for several representative charges and then analyze the outputs against expert annotations.
We select GPT-4o as the target LLM, as it achieves state-of-the-art performance on open-source legal benchmarks Fei et al. (2023); Li et al. (2024c) and also outperforms others in our pilot study (Appendix D), indicating a strong capacity to understand and apply legal knowledge. Following prior work Deng et al. (2023); Cui et al. (2024); Zhou et al. (2023), for each charge, we prompt GPT-4o with corresponding criminal articles (see prompt template in Appendix C).
We invite legal experts who passed the National Judicial Examination to analyze the LLM-generated FETs and identify two main issues:
(1) Inaccurate elements: LLMs may produce inaccurate FETs. For example, in Figure 1, for misappropriation of public funds, the LLM-generated Object is “the management order of public funds and the integrity of officials’ conduct”, missing the right to benefit from the use of public funds, which is necessary to identify this charge.
(2) Insufficient interpretive ability: LLMs fail to recognize when statutory language requires deeper interpretation. As shown in Figure 1, the model simply extracts “misappropriating public funds for personal use” to describe the Objective Aspect. However, this phrase is far too general for practice. In judicial interpretations [^2] National People’s Congress Standing Committee. Interpretation on Article 384, Paragraph 1 of the Criminal Law of the People’s Republic of China, adopted at the 27th Meeting of the Standing Committee of the 9th National People’s Congress on April 28, 2002., the term “for personal use” should be interpreted with three situations: (1) using public funds for oneself, relatives, or other individuals; (2) lending public funds to other entities in one’s own name; or (3) using public funds in the name of one’s organization for another entity to gain personal benefits.
@@ 5 Dataset Construction
The lack of accuracy and interpretation in the generated FETs undermines the reliability of legal reasoning tasks. To address this, we introduce an expert-annotated FET dataset that captures both formal legal definitions and practical interpretive nuances, supporting more trustworthy and adaptable legal AI systems.
To ensure both legal validity and interpretive clarity, we design a hierarchical annotation framework rooted in statutory sources and authoritative interpretive methods.
@@ 5.1 Hierarchical Legal Interpretation System
Figure 2: Hierarchical Legal Interpretation System based on legal source validity. The system consists of four annotation rounds, each using different interpretive methods based on different legal sources. Solid arrows indicate the primary method applied; dashed arrows represent supplementary use.
Given a specific charge, we ask five legal experts to annotate the FETs based on relevant legal materials like articles and cases. This annotation process is essentially an act of legal interpretation. Legal interpretation refers to the application of various methods to analyze and understand legal texts, to determine their meaning and application in specific legal contexts Barak (2005). In our task, it involves applying different interpretation methods to the different materials in order to analyze and define the connotation and extension of each of the FETs of a charge. In designing our annotation framework, we address the following two questions:
@@
(1) What sources are interpreted. Legal interpretation draws upon various legal sources with different levels of validity. In legal studies, these sources are categorized based on their legal validity into formal sources (which carry legal force in judgments) and informal sources (which serve as references without legal force) Pound (1925); Watson (1982); Pound (1932). Articles and judicial interpretations are considered formal sources, whereas case precedents and academic discourses are regarded as informal sources under the Chinese legal system Zhang and Zhou (2007). Accordingly, we organize legal sources by their level of validity, with the following order of priority: Article → Judicial Interpretations → Guiding Cases → Academic Discourses.
@@
(2) How the law is interpreted. When interpreting the above sources, different interpretation methods are required. These methods follow a hierarchical order Sutherland (1891); Kim and Division (2008); Eig (2014): Legal interpretation should begin with literal interpretation (interpreting the text based on its plain meaning). If the intended meaning cannot be clearly derived from the article alone, systematic interpretation (considering the article’s role within the legal system) and purposive interpretation (considering the legislative intent) should be applied. If ambiguity remains, historical interpretation (based on the legislative history), sociological interpretation (based on the article’s social function and consequences), and other interpretation methods may be used to further clarify the legal meaning. The specific definition of legal interpretation methods is in Appendix B.
We also consider the nature of each source. For example, Guiding Cases, as informal sources, do not define elements literally but instead supplement statutory interpretation through purposive, sociological, and other interpretive methods. The correspondence between interpretation methods and legal sources is illustrated by the orange arrows in Figure 2.
5.2 Annotation Process
As shown in the left part of Figure 2, our annotation process takes charges as input and outputs corresponding FETs, following a Hierarchical Legal Interpretation System to organize legal sources by validity and apply interpretation methods. The annotators are five experts, all of whom have passed the National Judicial Examination and are familiar with FET. The entire annotation process took 7 months and involved four rounds.
Stage One: Literal interpretation using the core article.
The interpretation of the FETs begins with the core article of each charge, which carries the highest legal validity, mainly through literal interpretation.
At this stage, annotators analyze the article’s subject–predicate–object structure to identify candidate FETs, mapping the subject to the Subject (who commits the crime), the predicate (verb phrase) to the Objective Aspect (the conduct carried out), the object to the Object (the legal interest infringed), and adverbial phrases to the Subjective Aspect (the offender’s mental state).
@@ 5.2 Annotation Process
For example, Article 263 of the Chinese Criminal Law, concerning robbery, states that “forcibly seizing public or private property through violence, coercion, or other means” describes the Objective Aspect. Since the article lacks an explicit subject, a general subject is assumed by default. The adverbs “violence” and “coercion” indicate an intentional act. This stage spans two months.
Stage Two: Systematic interpretation using related articles and judicial interpretations.
While Stage One relies on the primary article of each charge, literal interpretation often leaves FETs underspecified. Stage Two, therefore, applies systematic interpretation, situating underspecified terms within the broader legal framework. By considering the provision’s function in the Criminal Law and its links to related articles and judicial interpretations (broadly understood here to include SPC guiding opinions and other interpretative documents), annotators clarify the scope and meaning of the element.
For example, Article 263 of the Criminal Law does not specify whether “violence” must be directed only at persons or may also extend to property, nor does it make clear which borderline conduct should be excluded. Article 289 Congress (2017) provides that in mass “smashing, looting, and robbing,” the destruction or seizure of property by ringleaders may be punished as robbery, suggesting that violence may also cover acts against property. Conversely, the SPC’s 2016 Guiding Opinion on Robbery spc (2016) clarifies that if an offender uses only minor force to escape after theft, fraud, or snatching, and no injury above the statutory threshold results, such conduct is not deemed “violence” and does not requalify the offense as robbery.
Stage Three: Purposive and sociological interpretation using guiding cases.
Although the first two stages define the FETs based on articles and judicial interpretations, these sources remain abstract. In legal practice, courts also refer to Guiding Cases, designated by the Supreme People’s Court since 2011, to illustrate how legal articles are applied in concrete cases and interpreted in light of social purposes Chen et al. (2024).
At this stage, annotators refine FETs by incorporating specific case scenarios from Guiding Cases. Since the number of Guiding Cases is limited, for rarer charges, we also consult model cases from the People’s Court Case Database and Gazette cases333retrieved from PKULaw: https://www.pkulaw.com/case?way=topGuid. Annotators mainly apply purposive and sociological interpretation to examine how legal elements are concretized in the reasoning process of practical cases, considering both legislative intent and social context.
For example, in defining how “violence” in robbery operates in practice, annotators refer to a case involving molestation Ma (2021). The offender bound the victim to commit indecent acts and then took her phone. Under purposive and sociological interpretation, the ongoing molestation maintained the victim’s restrained state and thus constituted new violence. Such cases clarify that acts like molestation, though not physically injurious, can suppress resistance and therefore qualify as violence. This stage spans two months.
Stage Four: Diverse interpretations using academic discourses.
Although Stages One to Three refine both the abstract definitions and concrete scenarios of each element from various legal sources, certain elements still involve unresolved issues that rarely appear in practice and therefore lack clear judicial standards.
At this stage, annotators consult academic discourses and apply diverse interpretations such as comparative, purposive, and sociological interpretation. For elements where disagreement exists, they record both mainstream and minority views, providing concise annotations that explain the underlying legal reasoning.
For example, in defining “violence” in robbery, mainstream views in China, the former Soviet Union, North Korea, and Japan require that it endanger the victim’s life or health Zhang (2007b), while others argue that any force sufficient to subdue the victim should qualify Yang (2010). The annotations record both the dominant consensus and minority positions. This stage spans one month.
The four stages represent the main interpretive approaches, but they are not mutually exclusive. In practice, annotators often combine methods when clarifying a particular element. As shown in Figure 2, the dashed orange arrows mark cross-stage interactions where multiple interpretive methods operate in a complementary way.
5.3 Data Statistics
Metric LLMMean LLMMedian ExpertMean ExpertMedian